OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. 2 is a factor of n2 - n + 2

OpenStudy (anonymous):

@AnswerMyQuestions @texaschic101 @iGreen

OpenStudy (anonymous):

n^4 - 1 = (n^2 + 1)(n^2 - 1) = (n^2 + 1)(n - 1)(n + 1) is this correct

OpenStudy (texaschic101):

look at this... https://answers.yahoo.com/question/index?qid=20100110095352AACRmng

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

I didn't really do anything....I suppose you can thank the people on yahoo that answered this...lol

OpenStudy (anonymous):

lol but you helped :)